The firm GERANIUMS 'R' US produces POTTED GERANIUMS. The firm is considering whether to grow an essential input for that product: GERANIUM SEEDS. A total of 10,000 GERANIUM SEEDS are required each month. The firm currently buys the input for $0.15 per seed.

Growing GERANIUM SEEDS requires supplies and labor. The supplies would cost $475 per month and salaries would cost $375 per month.

Which of the following is TRUE?

a.

There IS  adequate information to perform relevant cost analysis and strategic cost analysis. Strategic cost analysis suggests the firm should make the product rather than buying it.

b.

It is $650 more  profitable for the firm to buy the input than to make it.

c.

It is $650 less profitable for the firm to buy the input than to make it.

d.

There is NOT adequate information to perform relevant cost analysis.
Answer #1

Answer is C

It is $650 less profitable for the firm to buy the input than to make it.

cost to produce the seeds = 475+375 =850

cost to buy = .15*1000 =1500

advantage to produce the seed over purchase = 1500-850 = 650